# Decoherence of Histories: Chaotic Versus Integrable Systems

### v2

@@ Abstract
Figure S6: Distribution of real ($f_{R}(\epsilon)$) and imaginary ($f_{I}(\epsilon)$) part of ${\epsilon}_{\boldsymbol{x},\boldsymbol{y}}$ in the XXZ model for $N=20$, $L=5$ and $T=T_{\text{eq}}$ for (i) chaotic; (ii) interacting integrable and (iii) free cases. $\sigma_{R,I}$ indicates the standard variance of $f_{R,I}(\epsilon)$.
@@
Figure S7: The expectation value $\langle B\rangle(t)$ of the coarse energy imbalance in the Ising model as a function of rescaled time for initial states $|\psi^{+}_{0}\rangle$ (solid line) and $|\psi^{-}_{0}\rangle$ (dashed line) for chaotic (i) and free (ii) cases. The dotted (dash-dotted) vertical line represents $T=T_{\text{neq}}$ ($T=T_{\text{eq}}$), where $T_{\text{neq}}$ indicates the time at which $\langle B\rangle(t)$ decays to $e^{-1}$ of its initial value and $T_{\text{eq}}=20T_{\text{neq}}$.
@@
We assume periodic boundary condition and set $J=h_{x}=1.0$. Two different values of $h_{z}$ are considered: (i) for $h_{z}=0.5$ (titled field) the system is chaotic; (ii) for $h_{z}=0.0$ (transverse field) the system is integrable and can be mapped to free fermions. A natural operator of interest here is an energy imbalance operator,
@@
$H^{R}=\sum_{\ell=\frac{N}{2}+1}^{N}(h_{x}\sigma_{x}^{\ell}+h_{z}\sigma_{z}^{\ell})+\sum_{\ell=\frac{N}{2}-1}^{N-1}J\sigma_{z}^{\ell}\sigma_{z}^{\ell+1}.$ (S3)
@@
$\Pi^{B}_{+}=\sum_{b_{k}\geq 0}|b_{k}\rangle\langle b_{k}|\quad\text{and}\quad\Pi^{B}_{-}=\sum_{b_{k}<0}|b_{k}\rangle\langle b_{k}|.$ (S4)
@@
Figure S8: Average $\overline{\epsilon}$ and maximum $\epsilon_{\text{max}}$ amount of coherence versus Hilbert space dimension $D$ in Ising model for the (i) chaotic (dark blue disks) and (ii) interacting integrable (medium blue diamonds) case. The dashed and dash-dotted line fit a scaling law of the form $D^{-\alpha}$ to (i) and (ii). The time step is $T=T_{\text{neq}}$ and the system sizes are $N=18,20,\dots,26$.
As a start, we plot the expectation value $\langle B\rangle(t)$ of the coarse energy imbalance as a function of time in Fig. S7. This is done for two different non-equilibrium initial states $|\psi_{0}^{\pm}\rangle$, where $|\psi_{0}^{\pm}\rangle$ is a Haar random state restricted to the subspace ±${\cal{H}}^{B}_{\pm}$. In the chaotic case (i) the system relaxes to its thermal equilibrium value $\langle B\rangle_{\text{eq}}=0$ with an equilibration time scale $\propto N^{2}$. In contrast, in the free case (ii) $\langle B\rangle(t)$ decays on a time scale $\propto N$ and fluctuations around $\langle B\rangle_{\text{eq}}$ remain visible for all times that we considered. Note, however, that—compared to the free case of the Heisenberg chain considered in the main text [cf. Fig. 1]—equilibration seems to work much better for the free case of the Ising model.
@@
The emergence of decoherence is investigated in Figs. S8 and S9 for Haar random initial states $|\psi_{0}\rangle$. We plot in double logarithmic scale $\overline{\epsilon}$ and $\epsilon_{\text{max}}$ versus the Hilbert space dimension $D$ for histories of lengths $L\in\{3,4,5\}$ for two different $T$: a nonequilibrium time scale $T_{\text{neq}}$ in Fig. S8 (identical to the dotted line in Fig. S7) and an equilibrium time scale $T_{\text{eq}}$ in Fig. S9 (identical to the dash-dotted line in Fig. S7). Finally, each data point in Figs. S8 and S9 is obtained by averaging $\overline{\epsilon}$ and $\epsilon_{\text{max}}$ over $2^{28-N}$ different realizations of $|\psi_{0}\rangle$.
Figure S10: Average $\overline{\epsilon}$ and maximum $\epsilon_{\text{max}}$ amount of coherence versus Hilbert space dimension $D$ for history length $L=5$ in the Ising model for chaotic (a,c) and non-interacting integrable (b,d) cases for time steps $T=0.5T_{\text{neq}},T_{\text{neq}},2T_{\text{neq}},3T_{\text{neq}},T_{\text{eq}}$ (from light to dark blue). The dashed line indicate the scaling $\sim D^{-1/2}$. The system size $N=18,20,\dots,26$.
Figure S11: Average $\overline{\epsilon}$ and maximum $\epsilon_{\text{max}}$ amount of coherence versus relative time step $T/T_{\text{neq}}$ for history length $L=5$ for the (i) chaotic (dark blue) and the (ii) non-interacting integrable (light blue) cases in the Ising model in double logarithmic scale. The system size is $N=22$
@@
For the chaotic case (i) we find a scaling law of the form $D^{-\alpha}$ (note that $D\propto 2^{N}$) with $\alpha\approx 0.25$ at the nonequilibrium time scale and with $\alpha\approx 0.5$ at the equilibrium time scale (with the same $\alpha$ for both $\overline{\epsilon}$ and $\epsilon_{\text{max}}$). This again indicates a robust exponential suppression (with respective to system size $N$) of coherences in chaotic systems. For the non-interacting integrable case (ii) we find smaller exponents $\alpha\approx 0.1$ at the nonequilibrium time scale, and $\alpha\approx 0$ at the equilibrium time scale.
The behaviour of decoherence as a function of $T$ is further studied in Fig. S10. For chaotic case (i) $\overline{\epsilon}$ and $\epsilon_{\text{max}}$ consistently decrease and the exponent grows to $\alpha\approx 0.5$ for increasing $T$. In contrast, for integrable case (ii), decoherence becomes (approximately) weaker for larger $T$, but it is again not as clear as for the example of the main text.
@@
Furthermore, similar to Fig. 4 in the main text, we plot $\overline{\epsilon}$ and $\epsilon_{\text{max}}$ versus $T$ for system size $N=22$ and history length $L=5$ in the Ising model in Fig. S11. Similar trends are observed: in chaotic case, both $\overline{\epsilon}$ and $\epsilon_{\text{max}}$ decrease with increasing $T$. In contrast, in the non-interacting integrable case, they increase on average (after a sudden drop at small $T$), accompanied by significant fluctuations.
To summarize, in the Ising model we find again a strong difference between the chaotic and integrable case in terms of the finite size scaling of $\overline{\epsilon}$ and $\epsilon_{\text{max}}$ and their behavior as a function of $T$. This is qualitatively in unison with what we found in the main text. However, we also found that the free case in the Ising model displays a much stronger form of decoherence compared to the free case in the Heisenberg model, showing that much care is required when one wants to generalize conclusions obtained for one specific integrable model.

### v3

@@ Abstract
Figure S5: Distribution of real ($f_{R}(\epsilon)$) and imaginary ($f_{I}(\epsilon)$) part of ${\epsilon}_{\boldsymbol{x},\boldsymbol{y}}$ in the XXZ model for $N=20$, $L=5$ and $T=T_{\text{eq}}$ for (i) chaotic; (ii) interacting integrable and (iii) free cases. $\sigma_{R,I}$ indicates the standard deviation of $f_{R,I}(\epsilon)$. Data from $2^{12}$ different initial Haar random states are taken into account.
S1.B. Results on spin density wave operator
In a translation invariant system as we considered here, we introduce the spin density wave operator with wave number $q\in\{-N/2,\dots,0,\dots,N/2-1\}$ (as in the main text we consider $N$ to be an even number)
$W_{q}^{0}=\sum_{\ell=1}^{N}\exp(-i\frac{2\pi q}{N}\ell)s_{z}^{\ell}.$ (S1)
Note that $W_{q=0}^{0}=\sum_{\ell=1}^{N}s_{z}^{\ell}$ is the total spin, which is conserved, and $W_{N-q}^{0}=(W_{q}^{0})^{\dagger}$. The spin imbalance operator studied in the main text can be expressed as
$A_{0}=\sum_{q=-N/2}^{N/2-1}c_{q}W^{0}_{q},$ (S2)
where
$c_{q}=\begin{cases}\frac{1}{N}{\displaystyle\left(i\cot\left(\frac{\pi q}{N}\right)-1\right)}&\text{odd\ }q\\ 0&\text{even\ }q\end{cases}.$ (S3)
Using $A_{0}=A_{0}^{\dagger}=\sum_{q=-N/2}^{N/2-1}c_{q}^{*}(W_{q}^{0})^{\dagger}$, the auto-correlation function of $A_{0}$ can be written as
$\langle A_{0}(t)A_{0}\rangle=\sum_{q,q^{\prime}=-N/2}^{N/2-1}c_{q^{\prime}}^{*}c_{q}\langle(W_{q^{\prime}}^{0}(t))^{\dagger}W_{q}^{0}\rangle,$ (S4)
where $\langle\bullet\rangle:=\frac{1}{D}\text{Tr}[\bullet]$. To further simply Eq. (S4), let us introduce the translation operator $\cal T$,
${\cal T}|\sigma_{1}\sigma_{2}\ldots\sigma_{L}\rangle=|\sigma_{L}\sigma_{1}\ldots\sigma_{L-1}\rangle,$ (S5)
where $|\sigma_{1}\sigma_{2}\ldots\sigma_{L}\rangle$ is the simultaneous basis of $s^{\ell}_{z}$, and $\sigma_{\ell}=\uparrow\downarrow$, indicating the spin up and down (along z-direction). It follows that ${\cal T}^{-1}s_{z}^{\ell}{\cal T}=s_{z}^{\ell-1}$ and
${\cal T}^{-1}W_{q}^{0}{\cal T}$ $=e^{-2\pi iq/N}W_{q}^{0},$
${\cal T}^{-1}(W_{q}^{0})^{\dagger}{\cal T}$ $=e^{2\pi iq/N}(W_{q}^{0})^{\dagger}.$ (S6)
Since the Hamiltonian is translational invariant ${\cal T}^{-1}H{\cal T}=H$, we find
$\begin{split}\langle W_{q^{\prime}}^{0}(t)^{\dagger}W_{q}^{0}\rangle&=\langle{(\cal T}^{-1}W_{q^{\prime}}^{0}{\cal T})(t)^{\dagger}{\cal T}^{-1}W_{q}^{0}{\cal T}\rangle\\ &=e^{2\pi i(q^{\prime}-q)/N}\langle(W_{q^{\prime}}^{0}(t))^{\dagger}W_{q}^{0}\rangle.\end{split}$ (S7)
This equation can only be satisfied if $\langle W_{q^{\prime}}^{0}(t)^{\dagger}W_{q}^{0}\rangle=0$ for $q\neq q^{\prime}$ or $q=q^{\prime}$. Thus, Eq. (S4) becomes
$\langle A_{0}(t)A_{0}\rangle=\sum_{q=-N/2}^{N/2-1}|c_{q}|^{2}\langle W_{q}^{0}(t)^{\dagger}W_{q}^{0}\rangle,$ (S8)
Assuming diffusive spin transport, which has been confirmed numerically for our cases $\Delta_{1}=1.5,\Delta_{2}=0.5$ and $\Delta_{1}=1.5,\Delta_{2}=0$ (see Ref. [68] and references therein), we have
$\langle W_{q}^{0}(t)^{\dagger}W_{q}^{0}\rangle\sim e^{-(\frac{2\pi q}{N})^{2}D_{c}t},\ \text{for}\ \frac{|q|}{N}\rightarrow 0,$ (S9)
where $D_{c}$ denotes the diffusion constant. For $N\rightarrow\infty$, Eq. (S3) becomes
$|c_{q}|\simeq\begin{cases}{\displaystyle\frac{1}{\pi q}}&\text{odd}\ q,\ \frac{|q|}{N}\rightarrow 0\\ 0&\text{otherwise}\end{cases}.$ (S10)
As a result, the auto-correlation function of the spin imbalance operator $\langle A_{0}(t)A_{0}\rangle$ is dominated by the auto-correlation function of the density wave operator $\langle W^{0}_{q}(t)^{\dagger}W^{0}_{q}\rangle$ at small $q$, especially the late time behavior.
Figure S6: $\langle\widetilde{W}_{q=1}\rangle(t)$ as a function of rescaled time for initial states $|\psi^{+}_{0}\rangle$ (solid line) and $|\psi^{-}_{0}\rangle$ (dashed line) for the chaotic (i), interacting integrable (ii) and non-interacting integrable (iii) cases. The dotted (dash-dotted) vertical line represents $T=T_{\text{neq}}$ ($T=T_{\text{eq}}$), where $T_{\text{neq}}$ indicates the time at which $\langle\widetilde{W}_{q=1}\rangle(t)$ decays to $e^{-1}$ of its initial value and $T_{\text{eq}}=20T_{\text{neq}}$. Specifically, we obtain $T_{\text{neq}}$ from the data of the largest system size $(N_{\text{max}}=28)$. Then, for smaller sizes $N$ we set $T_{\text{neq}}(N)=(N/N_{\text{max}})^{p}T_{\text{neq}}(N_{\text{max}})$ with $p=2$ for (i) and (ii) and $p=1$ for (iii).
Figure S7: Average $\overline{\epsilon}$ and maximum $\epsilon_{\text{max}}$ amount of coherence versus Hilbert space dimension $D$ for the (i) chaotic (dark blue disks), (ii) interacting integrable (medium blue diamonds) and (iii) free (light blue triangles) case for $L\in\{3,4,5\}$ of the coarse spin density wave $\widetilde{W}_{q=1}$. The dashed and dash-dotted line fit a scaling law of the form $D^{-\alpha}$ to (i) and (ii). The time step is $T=T_{\text{neq}}$ and the system sizes are $N=18,20,\dots,28$. Note the double-logarithmic scale.
Figure S8: Identical to Fig. S7 except for time steps $T=T_{\text{eq}}$.
Figure S9: Decoherence as a function of the time step $T=t_{k}-t_{k-1}$ for history length $L=5$ of the coarse spin density wave $\widetilde{W}_{q=1}$. (a,b) $\overline{\epsilon},\epsilon_{\text{max}}$ for $N=26$ on a double logarithmic scale. (c,d) Scaling exponent $\alpha$ for $\overline{\epsilon},\epsilon_{\text{max}}$ obtained for system sizes $N=18,20,\ldots,28$ and as a mean over $2^{30-N}$ realizations of a Haar random $|\psi_{0}\rangle$. Red dashed line in (c,d) indicates $\alpha=0.5$. Black dotted (dash-dotted) line indicates $T=T_{\text{neq}}(T=T_{\text{eq}})$.
In the following, we first consider the DF corresponding to the density wave operator $W^{0}_{q}$ for the slowest (non-trivial) mode $q=1$. For simplicity, we also consider the Hermitian version of it,
$\widetilde{W}_{q}^{0}=\sum_{\ell=1}^{N-1}\cos(\frac{2\pi q}{N}\ell)s_{z}^{\ell}.$ (S11)
The coarse grained operator can be constructed as, $\widetilde{W}_{q}=\Pi^{W}_{+}-\Pi^{W}_{-}$ with projectors
$\Pi^{W}_{+}=\sum_{w_{k}>0}|w_{k}\rangle\langle w_{k}|\quad\text{and}\quad\Pi^{W}_{-}=\sum_{w_{k}\leq 0}|w_{k}\rangle\langle w_{k}|,$ (S12)
where $w_{k}$ and $|w_{k}\rangle$ denote the eigenvalue and eigenvectors of $\widetilde{W}^{0}_{q=1}$, respectively. For all system sizes considered here, we focus on the subspace with $S_{z}=0$.
Results for $\widetilde{W}_{q=1}$ are shown in Figs. S6S7S8 and S9, which are analogous to Figs. 1, 2, 3 and 4 of the main text. We consistently observe the same qualitative behavior as in the main text with small quantitative differences. In particular, we again observe strong differences between the (i) chaotic, (ii) interacting integrable and (iii) free case. This clearly indicates that the results of the main text are general.
We remark that this nicely ties together the ideas of van Kampen (that slow observables can be described by a classical stochastic process) and of Gell-Mann, Hartle, Halliwell and others (that quasi-conserved hydrodynamic modes give rise to decoherent histories). In fact, the spin imbalance operator considered in the main text evolves on a slow time scale (compared to the shortest microscopic evolution time scale determined by the quantum speed limit), which is echoed here as its long time behaviour is dominated by large wavelength ($q/N\approx 0$ or “quasi-conserved”) hydrodynamic modes.
Figure S10: Similar to Fig. S6, but for $\widetilde{W}_{q=N/2}$. The dotted (dash-dotted) vertical line represents $T=T_{\text{neq}}$ ($T=T_{\text{eq}}$), where $T_{\text{neq}}$ indicates the time at which $\langle\widetilde{W}_{q=N/2}\rangle(t)$ (for largest system size considered $N_{\text{max}}=28$) decays to $e^{-1}$ of its initial value and $T_{\text{eq}}=20T_{\text{neq}}$.
Figure S11: Similar to Fig. S7, but for $\widetilde{W}_{q=N/2}$.
Furthermore, we also investigate the fast mode $q=N/2$. In contrast to the case $q=1$, $\langle\widetilde{W}_{q=N/2}\rangle(t)$ decays on a much shorter time scale (Fig. S10), which is approximately independent of the system size, particularly for the chaotic and interacting integrable case. In other words, $\widetilde{W}_{q=N/2}$ is a fast observable. The emergence of consistent decoherent histories with respect to fast observables is a subtle topic, here we restrict ourselves to reporting the numerical results without delving into detailed discussions.
Figure S12: Similar to Fig. S8, but for $\widetilde{W}_{q=N/2}$.
Figure S13: Similar to Fig. S9, but for $\widetilde{W}_{q=N/2}$.
We choose $S_{z}=0$ for system size $N=4k+2$ and $S_{z}=1$ for $N=4k$, where $k\in\mathbb{N}$. Results are shown in Figs. S10S11S12 and S13, which are analogous to Figs. S6S7S8 and S9. As expected, notably different results are observed compared to the slow mode density wave operator $\widetilde{W}_{q=1}$, especially for the chaotic and interacting integrable case:
  1. 1.
    At the non-equilibrium time scale $T_{\text{neq}}$ (Fig. S11), we observe $\alpha\approx 0$ even in the chaotic case indicating a possible sub-exponential or power-law suppression (with respect to $N$) of coherences. Surprisingly, decoherence is even slightly weaker compared to the interacting integrable case.
  2. 2.
    At the equilibrium time scale $T_{\text{eq}}$ (Fig. S12), decoherence is recovered in the chaotic case ($\alpha\approx 0.5$), and a similar behavior appears to hold in the interacting integrable case with a smaller $\alpha\approx 0.3$.
  3. 3.
    The overall behavior of the scaling exponent $\alpha$ as a function of $T$ (Fig. S13) is similar to the slow mode except for two differences. First, the onset of decoherence happens later for all cases. Second, the scaling exponent in case (ii) remains finite ($\alpha\approx 0.25$) and does not decay to zero.
To conclude, we found a noticeably different behaviour of the fast mode compared to the three slow cases (the spin imbalance in the main text, the slow $q=1$ mode above and the example in the next section). In particular, our results show that coarseness alone is not sufficient to guarantee decoherence for nonequilibrium time scales, even in a chaotic system. However, in unison with the three slow cases we also found here a distinct behavior for the (i) chaotic, (ii) interacting integrable, and (iii) non-interacting integrable cases for large times.
@@
We assume periodic boundary condition and set $J=h_{x}=1.0$. Two different values of $h_{z}$ are considered: (i) for $h_{z}=0.5$ (titled field) the system is chaotic; (iii) for $h_{z}=0.0$ (transverse field) the system is integrable and can be mapped to free fermions. Note that the system is not known to have an interacting integrable case (ii). A natural operator of interest here is an energy imbalance operator,
@@
$H^{R}=\sum_{\ell=\frac{N}{2}+1}^{N}(h_{x}\sigma_{x}^{\ell}+h_{z}\sigma_{z}^{\ell})+\sum_{\ell=\frac{N}{2}+1}^{N-1}J\sigma_{z}^{\ell}\sigma_{z}^{\ell+1}.$ (S15)
@@
$\Pi^{B}_{+}=\sum_{b_{k}>0}|b_{k}\rangle\langle b_{k}|\quad\text{and}\quad\Pi^{B}_{-}=\sum_{b_{k}\leq 0}|b_{k}\rangle\langle b_{k}|.$ (S16)
@@
Figure S14: The expectation value $\langle B\rangle(t)$ of the coarse energy imbalance in the Ising model as a function of rescaled time for initial states $|\psi^{+}_{0}\rangle$ (solid line) and $|\psi^{-}_{0}\rangle$ (dashed line) for chaotic (i) and free (iii) cases. The dotted (dash-dotted) vertical line represents $T=T_{\text{neq}}$ ($T=T_{\text{eq}}$), where $T_{\text{neq}}$ indicates the time at which $\langle B\rangle(t)$ decays to $e^{-1}$ of its initial value and $T_{\text{eq}}=20T_{\text{neq}}$. Specifically, we obtain $T_{\text{neq}}$ from the data of the largest system size $(N_{\text{max}}=26)$. Then, for smaller sizes $N$ we set $T_{\text{neq}}(N)=(N/N_{\text{max}})^{p}T_{\text{neq}}(N_{\text{max}})$ with $p=2$ for (i) and $p=1$ for (iii). The free case is denoted as (iii) to remain consistent with the notation used in the main text.
As a start, we plot the expectation value $\langle B\rangle(t)$ of the coarse energy imbalance as a function of time in Fig. S14. This is done for two different non-equilibrium initial states $|\psi_{0}^{\pm}\rangle$, where $|\psi_{0}^{\pm}\rangle$ is a Haar random state restricted to the subspace ±${\cal{H}}^{B}_{\pm}$. In the chaotic case (i) the system relaxes to its thermal equilibrium value $\langle B\rangle_{\text{eq}}=0$ with an equilibration time scale $\propto N^{2}$. In contrast, in the free case (iii) $\langle B\rangle(t)$ decays on a time scale $\propto N$ and fluctuations around $\langle B\rangle_{\text{eq}}$ remain visible for all times that we considered. Note, however, that—compared to the free case of the Heisenberg chain considered in the main text [cf. Fig. 1]—equilibration seems to work much better for the free case of the Ising model.
Figure S15: Average $\overline{\epsilon}$ and maximum $\epsilon_{\text{max}}$ amount of coherence versus Hilbert space dimension $D$ of the coarse energy imbalance in Ising model for the (i) chaotic (dark blue disks) and (ii) interacting integrable (medium blue diamonds) case. The dashed and dash-dotted line fit a scaling law of the form $D^{-\alpha}$ to (i) and (ii). The time step is $T=T_{\text{neq}}$ and the system sizes are $N=18,20,\dots,26$.
@@
The emergence of decoherence is investigated in Figs. S15 and S16 for Haar random initial states $|\psi_{0}\rangle$. We plot in double logarithmic scale $\overline{\epsilon}$ and $\epsilon_{\text{max}}$ versus the Hilbert space dimension $D$ for histories of lengths $L\in\{3,4,5\}$ for two different $T$: a nonequilibrium time scale $T_{\text{neq}}$ in Fig. S15 (identical to the dotted line in Fig. S14) and an equilibrium time scale $T_{\text{eq}}$ in Fig. S16 (identical to the dash-dotted line in Fig. S14). Finally, each data point in Figs. S15 and S16 is obtained by averaging $\overline{\epsilon}$ and $\epsilon_{\text{max}}$ over $2^{30-N}$ different realizations of $|\psi_{0}\rangle$.
Figure S17: Decoherence as a function of the time step $T=t_{n}-t_{n-1}$ for history length $L=5$ of the energy imbalance in the Ising model for the (i) chaotic (dark blue) and the (iii) free (light blue) cases. (a,b) $\overline{\epsilon},\epsilon_{\text{max}}$ for $N=24$ on a double logarithmic scale. (c,d) Scaling exponent $\alpha$ for $\overline{\epsilon},\epsilon_{\text{max}}$ obtained for system sizes $N=18,20,\ldots,26$ and as a mean over $2^{28-N}$ realizations of a Haar random $|\psi_{0}\rangle$. Black dotted (dash-dotted) line indicates $T=T_{\text{neq}}(T=T_{\text{eq}})$. Red dashed line in (c,d) indicates $\alpha=0.5$.
@@
For the chaotic case (i) we find a scaling law of the form $D^{-\alpha}$ (note that $D\propto 2^{N}$) with $\alpha\approx 0.25$ at the nonequilibrium time scale and with $\alpha\approx 0.5$ at the equilibrium time scale (with the same $\alpha$ for both $\overline{\epsilon}$ and $\epsilon_{\text{max}}$). This again indicates a robust exponential suppression (with respective to system size $N$) of coherences in chaotic systems. For the free case (iii) we find smaller exponents $\alpha\approx 0.1$ at the nonequilibrium time scale, and $\alpha\approx 0$ at the equilibrium time scale.
@@
Furthermore, similar to Fig. 4 in the main text, we plot $\overline{\epsilon}$ and $\epsilon_{\text{max}}$ for system size $N=24$ as well as the scaling exponent $\alpha(\overline{\epsilon})$ ($\alpha({\epsilon}_{\text{max}})$) versus $T$ for history length $L=5$ in the Ising model in Fig. S17. Similar trends are observed: in chaotic case, both $\overline{\epsilon}$ and $\epsilon_{\text{max}}$ decrease with increasing $T$. In contrast, in the free case, they increase on average (after a sudden drop at small $T$), accompanied by significant fluctuations. The general behavior of the scaling exponent $\alpha(\overline{\epsilon})$ ($\alpha({\epsilon}_{\text{nax}})$), especially for time scale $T\geq T_{\text{neq}}$, is also similar to the results shown in Fig. 4. This cements the idea that the decoherence behavior of slow observables is in general characterized by common qualitative features that clearly distinguish between integrable and non-integrable dynamics.